# The Tianlai-WIYN North Celestial Cap Redshift Survey  [^1] Template AASArticle with Examples: v6.3.1Released on March, 1st, 2021

### v1

@@ Abstract
We present the results of a small, low redshift spectroscopic survey of galaxies within $3^{\circ}$ of the North Celestial Pole (NCP) selected using V-band photometry obtained from the North Celestial Cap Survey (NCCS) (Gorbikov_2014). The purpose of the current survey is to create a redshift space template for 21 cm emission from neutral hydrogen with which to correlate radio line intensity observations by the Tianlai dish and cylinder interferometers. A total of 898 redshifts were obtained from the 2102 extended objects in the NCCS with $m_{V}<19$ in the survey area. After accounting for extinction, the survey geometry and selection effects, the number density and clustering pattern of galaxies in the redshift catalog are consistent with other low redshift surveys. We were also able to identify 11 galaxy cluster candidates from this redshift catalog.
@@ I Introduction
In this paper we present the results of a spectroscopic optical galaxy redshift survey of the North Celestial Cap (NCC). This survey was done in support of the Tianlai Pathfinder (2021MNRAS.506.3455W; 2020SCPMA..6329862L), a radio telescope array being used to develop the 21 cm intensity mapping technique (2008PhRvL.100i1303C; 2008PhRvL.100p1301L; 2020PASP..132f2001L). For decades, the 21 cm radio emission line has been used to construct redshift catalogs of individual galaxies with depths limited by the sensitivity to individual galaxies (2019MNRAS.486.5124O). 21 cm intensity mapping can function at greater depths where individual galaxies are unresolved, determining the large scale spatial distribution of the neutral hydrogen gas. Indeed, 21 cm intensity mapping observations have set upper limits on the the spatial distribution of hydrogen gas at redshifts spanning the Epoch of Reionization and Cosmic Dawn. They may eventually measure the distribution of hydrogen gas during the cosmic Dark Ages, before galaxies formed, where most of the gas is neutral hydrogen. Observations in all redshift ranges suffer from common challenges, the most significant being radio foregrounds orders of magnitude brighter than the hydrogen signal. Until recently, measurements by individual instruments have resulted only in upper limits on the hydrogen signal, with the exception of measurements at $z=0.33$ and $z=0.44$ using the MeerKAT (Paul2023) interferometer and at $z\sim 1$ with the CHIME interferometer (CHIME2025). In addition, successful detections have been achieved by cross-correlations with optical surveys. The CHIME interferometer has also successfully detected HI emission by cross-correlating with large-scale structure in the eBOSS survey (2023ApJ...947...16A; 2024ApJ...963...23A) as has the MeerKAT interferometer (operated as multiple single dishes) (2023MNRAS.518.6262C) with the Wiggle-Z galaxy redshift survey. Single dish instruments (e.g. Wolz_2021 with GBT data, Anderson_2018 with Parkes data) have done so as well. Observing cross-correlation signals at levels on par with detailed simulations validates the understanding both of systematic effects and the analysis pipelines. So far, in redshift ranges where galaxy redshift surveys exist, 21 cm intensity mapping results are far from being competitive with them. This is largely due to a variety of systematic effects which it is hoped can be ameliorated.
@@
Traditional 21 cm galaxy redshift surveys operate in a regime where individual galaxies can be differentiated. At somewhat larger $z$’s there is an intermediate (“hybrid”) regime where one can identify individual large galaxies or groups of galaxies but where most of the 21 cm signal comes from unresolved galaxies which can be studied using 21 cm intensity mapping. This is the regime in which the Tianlai Dish Array (TDA) low redshift survey will operate. It is expected to detect the 2121\,cm signal, including a few HI “clumps” (2022MNRAS.517.4637P). Currently, the TDA correlator has an instantaneous bandwidth of $\sim 100\,\mathrm{MHz}$ that can be tuned within the range from 4001430400-1430\,MHz (2.55>z>0.012.55>z>-0.01). The nearest volume that can be surveyed with a single tuning of the array includes $0\lesssim z\lesssim 0.07$. Just as other 21 cm intensity mapping surveys at higher redshift have benefited from cross-correlation with galaxy redshift surveys, having a spectroscopic optical galaxy redshift survey overlapping the TDA survey volume can help to disentangle signal from systematics. This is the purpose of the current survey.
@@
One systematic uncertainty is the shape of the radio beam, particularly the sidelobes. Another is electromagnetic interactions between different elements of the array (Kern2019; Kern2020). Both of these are sensitive to the array’s configuration. To minimize the complexity of these uncertainties, arrays designed for 21 cm intensity mapping generally operate as transit telescopes pointing at a fixed declination and scanning the sky over right ascension as the Earth rotates. Since there is a much larger solid angle at low declination (near the celestial equator) than at high declination (near the NCP), the telescope will spend a smaller fraction of time observing any given patch of the sky at low declination than at high. This results in poorer brightness temperature (or intensity) sensitivity at low declination for a given observing time. At the other extreme, if one points at the Celestial Poles (North NCP or South SCP) then one is always observing the same patch of the sky, which leads to far better sensitivity. For this reason, most of the TDA observations to date (near $z\sim 1$) were performed with the array pointing directly at the NCP. We chose the same observing strategy for the low redshift survey. At the lowest $z$’s the observational wavelength is $\lambda\sim 21\,\mathrm{cm}$ and the TDA angular resolution is $\sim 2.3^{\circ}$. Thus, to complement the low-$z$ TDA survey, one wants an optical galaxy spectroscopic redshift survey of the $z\leq 0.07$ volume within $\sim 2^{\circ}$ of the NCP. We refer to our survey area around the NCP as the North Celestial Cap or NCC and use tNCCSz as an acronym for our redshift survey.
@@
The Tianlai collaboration carried out the spectroscopic survey prescribed above because no such survey existed. (Fortunately, the NCCS photometric survey (Gorbikov_2014) had already been performed for other reasons.) There are many reasons why this survey has not been done before:
@@
Figure 1: Left panel: Planck V-band extinction map showing $A_{v}$ (in magnitudes) in a 6 radius region around the NCP. We have used the Planck dust extinction map (2016A&A...586A.132P) to model the Galactic extinction. Note that the tNCCSz survey covers the central region to a radius of 3. One can notice that this area is subject to significant absorption by dust, reaching 1.7 magnitude in some directions covered by the survey. Right panel: Distribution of targets and associated tiling map for the tNCCSz survey. Color-coded by magnitude, the map highlights galaxies in a specific magnitude range for redshift analysis. The circular red outlines represent the 56 observational tiles, each corresponding to a unique pointing of the WIYN telescope to cover the declination range above $87^{\circ}$ and parts of the 86.6 to 87.0 range. The blued dashed circle represents the full width at half maximum of the primary beams of the TDA antennas at 1380 MHz ($z\sim 0.035$).
@@ II The Survey
To conduct an optical galaxy redshift survey, one must first identify the angular positions of the galaxies using a photometric survey and then perform spectroscopic observations. The photometry for our study was previously done in the North Celestial Cap Survey (NCCS Gorbikov_2014). The NCCS identifies approximately $4\times 10^{6}$ objects within $10^{\circ}$ of the NCP using V-R-I photometry with magnitude limits $m_{V}<20.3$, $m_{R}<21.0$, and $m_{I}<19.2$. Some of these objects are stars, and some galaxies, particularly at the magnitude limit, are likely beyond our primary interest in redshifts with $z<=0.07$. Thus, only a fraction of the NCCS survey is useful for our purposes. The catalog also includes a PESS (point-extended-source-separation) index indicating whether an object is point-like (PESS <=<= 1, most likely a star) or extended (PESS = 2 or 3, most likely a galaxy). Selecting spectroscopic targets involved cutting on declination and magnitude and excluding point-like objects.
@@ II.2 Observations
Between February 2020 and May 2023, we obtained 15 nights of gray and dark time on the WIYN 3.5 m telescope to use the Hydra multifiber positioner and bench spectrograph (1995SPIE.2476...56B) to survey the NCC. We used the bench spectrograph with the blue-light optimized fibers and the 400@4.2 (WIYN Hydra and Bench spectrograph users manual - v5) grating. These choices resulted in wavelength coverage from below 3800 Å to above 7600 Å. However, our sensitivity below about 4600 Å was typically insufficient to identify the common absorption features in galaxy spectra, and the $\textrm{H}_{\alpha}$ emission line could be confused with prominent sky lines above 7200 Å.
@@ III.3.2 Semi-automatic Method
We chose to use two (for robustness) galaxy spectra templates with emission features and two with absorption features. Emission line galaxy spectra have been chosen from the SDSS DR5 spectral cross-correlation templates [^2] https://classic.sdss.org/dr5/algorithms/spectemplates/. Specifically, we used templates with IDs 24 and 25 in our processing. For absorption line galaxies, on the one hand, we used the NGC 4125 spectrum listed by Brown_2014 in their representative subsample. We also used template no. 41 in the collection used in the AUTOZ code developed within the GAMA collaboration (Baldry_2014)333code and data can be found in I. Baldry’s website: https://www.astro.ljmu.ac.uk/ ikb/research/. Finally, two redshift measurements have been derived using template ID 30 in the SDSS DR5 collection, corresponding to objects with broad emission lines (like QSOs or AGN). Prior to their use in redshift determination, all templates are pre-processed by fitting and subtracting a low order polynomial.
@@ IV.1.1 A posteriori check of target selection with Gaia
In order to estimate the fraction of proposed targets that might actually be stars, rather than galaxies, we matched the NCCS catalog with the Gaia DR3 catalog (gaia_collaboration_gaia_2023). As this catalog was released well after the beginning of our survey, this is only an a posteriori check. First, we looked at the distribution of the PESS flag for NCCS objects with $\delta\geq 86.6^{\circ}$ and $13<m_{V}<18.995$ - essentially the same as the criteria used for our target selection - shown on Figure 6. Within this sample, 1714 and 669 objects have PESS flag equal to 2 and 3, respectively. Of these, 1046 objects with PESS = 2, and 222 objects with PESS = 3, are closer than 2 arcsec from an object in Gaia DR3 (a “match” in the following). In contrast, more than 99.8% of NCCS objects with PESS $\leq$ 1 (point-like) have a match in Gaia DR3.
@@ IV.1.3 Comparison with the 2MASS Redshift Survey
The 2MASS redshift survey catalog (huchra_2mass_2012) contains a number of galaxies that overlap our survey area. As a means for testing the completeness of the tNCCSz, we searched the 2MASS catalog for galaxies with $\delta\geq 86.5^{\circ}$ and $m_{K}>13.5$. We found 26 such galaxies and
our tNCCSz survey obtained spectra for 14 of them. The breakdown of the reasons the remaining 12 were not in our sample is summarized in Table 2. The two targets that were “Not in the NCCS Catalog” were detected in V-band by NCCS but without corresponding measurements in R or I bands. These objects appear in the full NCCS catalog but not in the “cleaned” catalog that we used for target selection. Somewhat surprisingly, 10 galaxies were not included in our target list because the NCCS catalog PESS score indicated that they were more likely stars. This result appears to be inconsistent with our understanding that the NCCS catalog separation between stars and galaxies is $92\%$ accurate, based on a comparison of NCCS with SDSS (Gorbikov_2014).
@@ IV.2 Selection Function
Using the above efficiency functions, we are able to reproduce the magnitude distribution of our final list of galaxies with redshifts, the NCCS catalog itself, and the selected targets for this survey. To do so, we start from the galaxy absolute luminosity distribution function, and we take into account the absorption by the Galactic dust, with its variations over the $\sim 35\,\mathrm{deg}^{2}$ survey area (Figure 1, right panel). The galaxy luminosity distribution is modeled as a Schechter function (see Eq. 7) with the parameters given by 2009A&A...508.1217Z. More precisely, we have used a characteristic magnitude $M^{*}=-20.5$, normalisation $\Phi^{*}=0.0063\mathrm{Mpc}^{-3}$ and a faint end slope $\alpha=-1.07$ as the luminosity function parameters. More details can be found in paragraph V.3.1.
@@ IV.3 Catalog Description
The tNCCSz redshift catalog contains a total of 23 columns and 1874 rows, with each row corresponding to an NCCS (Gorbikov_2014) source, with the associated redshift information extracted from the current survey’s observations. Eleven columns correspond to information in the original NCCS catalog, and an additional 12 columns contain redshift information derived from this survey. The full table in electronic form will be available from the the Centre de Données astronomiques de Strasbourg (CDS) [^5] CDS, Strasbourg astronomical Data Center: https://cds.unistra.fr.
@@ V Clustering analysis
One way to validate tNCCSz is to analyze the redshift space clustering of our galaxies and compare this with other surveys since the clustering statistics in different cosmological volumes are expected to be the same. Due to its small solid angle tNCCSz can add little new to our knowledge of galaxy clustering, but this is not why this survey was made. The comparison we make is with the Sloan Digital Sky Survey (SDSS). SDSS is chosen because of its similar depth, excellent control of systematics, and its very large area, which provides a large statistical sample. More specifically, we compare tNCCSz with mock tNCCSz catalogs derived from the SDSS spectroscopic catalog (2017AJ....154...28B). The procedure for generating these mocks is described in § V.1.

### v2

@@ Abstract
We present the results of a small, low redshift spectroscopic survey of galaxies within $3^{\circ}$ of the North Celestial Pole (NCP) selected using V-band photometry obtained from the North Celestial Cap Survey (NCCS) (Gorbikov and Brosch, 2014). The purpose of the current survey is to create a redshift space template for 21 cm emission from neutral hydrogen with which to correlate radio line intensity observations by the Tianlai dish and cylinder interferometers. A total of 898 redshifts were obtained from the 2102 extended objects in the NCCS with $m_{V}<19$ in the survey area. After accounting for extinction, the survey geometry and selection effects, the number density and clustering pattern of galaxies in the redshift catalog are consistent with other low redshift surveys. We were also able to identify 11 galaxy cluster candidates from this redshift catalog.
@@ I Introduction
In this paper we present the results of a spectroscopic optical galaxy redshift survey of the North Celestial Cap (NCC). This survey was done in support of the Tianlai Pathfinder (Wu et al., 2021; Li et al., 2020), a radio telescope array being used to develop the 21 cm intensity mapping technique (Chang et al., 2008; Loeb and Wyithe, 2008; Liu and Shaw, 2020). For decades, the 21 cm radio emission line has been used to construct redshift catalogs of individual galaxies with depths limited by the sensitivity to individual galaxies (Obuljen et al., 2019). 21 cm intensity mapping can function at greater depths where individual galaxies are unresolved, determining the large scale spatial distribution of the neutral hydrogen gas. Indeed, 21 cm intensity mapping observations have set upper limits on the the spatial distribution of hydrogen gas at redshifts spanning the Epoch of Reionization and Cosmic Dawn. They may eventually measure the distribution of hydrogen gas during the cosmic Dark Ages, before galaxies formed, where most of the gas is neutral hydrogen. Observations in all redshift ranges suffer from common challenges, the most significant being radio foregrounds orders of magnitude brighter than the hydrogen signal. Until recently, measurements by individual instruments have resulted only in upper limits on the hydrogen signal, with the exception of measurements at $z=0.33$ and $z=0.44$ using the MeerKAT (Paul et al., 2023) interferometer and at $z\sim 1$ with the CHIME interferometer (Amiri et al., 2025). In addition, successful detections have been achieved by cross-correlations with optical surveys. The CHIME interferometer has also successfully detected HI emission by cross-correlating with large-scale structure in the eBOSS survey (Amiri et al., 2023, 2024) as has the MeerKAT interferometer (operated as multiple single dishes) (Cunnington et al., 2023) with the Wiggle-Z galaxy redshift survey. Single dish instruments (e.g. Wolz et al. (2021) with GBT data, Anderson et al. (2018) with Parkes data) have done so as well. Observing cross-correlation signals at levels on par with detailed simulations validates the understanding both of systematic effects and the analysis pipelines. So far, in redshift ranges where galaxy redshift surveys exist, 21 cm intensity mapping results are far from being competitive with them. This is largely due to a variety of systematic effects which it is hoped can be ameliorated.
@@
Traditional 21 cm galaxy redshift surveys operate in a regime where individual galaxies can be differentiated. At somewhat larger $z$’s there is an intermediate (“hybrid”) regime where one can identify individual large galaxies or groups of galaxies but where most of the 21 cm signal comes from unresolved galaxies which can be studied using 21 cm intensity mapping. This is the regime in which the Tianlai Dish Array (TDA) low redshift survey will operate. It is expected to detect the 2121\,cm signal, including a few HI “clumps” (Perdereau et al., 2022). Currently, the TDA correlator has an instantaneous bandwidth of $\sim 100\,\mathrm{MHz}$ that can be tuned within the range from 4001430400-1430\,MHz (2.55>z>0.012.55>z>-0.01). The nearest volume that can be surveyed with a single tuning of the array includes $0\lesssim z\lesssim 0.07$. Just as other 21 cm intensity mapping surveys at higher redshift have benefited from cross-correlation with galaxy redshift surveys, having a spectroscopic optical galaxy redshift survey overlapping the TDA survey volume can help to disentangle signal from systematics. This is the purpose of the current survey.
@@
One systematic uncertainty is the shape of the radio beam, particularly the sidelobes. Another is electromagnetic interactions between different elements of the array (Kern et al., 2019, 2020). Both of these are sensitive to the array’s configuration. To minimize the complexity of these uncertainties, arrays designed for 21 cm intensity mapping generally operate as transit telescopes pointing at a fixed declination and scanning the sky over right ascension as the Earth rotates. Since there is a much larger solid angle at low declination (near the celestial equator) than at high declination (near the NCP), the telescope will spend a smaller fraction of time observing any given patch of the sky at low declination than at high. This results in poorer brightness temperature (or intensity) sensitivity at low declination for a given observing time. At the other extreme, if one points at the Celestial Poles (North NCP or South SCP) then one is always observing the same patch of the sky, which leads to far better sensitivity. For this reason, most of the TDA observations to date (near $z\sim 1$) were performed with the array pointing directly at the NCP. We chose the same observing strategy for the low redshift survey. At the lowest $z$’s the observational wavelength is $\lambda\sim 21\,\mathrm{cm}$ and the TDA angular resolution is $\sim 2.3^{\circ}$. Thus, to complement the low-$z$ TDA survey, one wants an optical galaxy spectroscopic redshift survey of the $z\leq 0.07$ volume within $\sim 2^{\circ}$ of the NCP. We refer to our survey area around the NCP as the North Celestial Cap or NCC and use tNCCSz as an acronym for our redshift survey.
@@
The Tianlai collaboration carried out the spectroscopic survey prescribed above because no such survey existed. (Fortunately, the NCCS photometric survey (Gorbikov and Brosch, 2014) had already been performed for other reasons.) There are many reasons why this survey has not been done before:
@@
Figure 1: Left panel: Planck V-band extinction map showing $A_{v}$ (in magnitudes) in a 6 radius region around the NCP. We have used the Planck dust extinction map (Planck Collaboration et al., 2016) to model the Galactic extinction. Note that the tNCCSz survey covers the central region to a radius of 3. One can notice that this area is subject to significant absorption by dust, reaching 1.7 magnitude in some directions covered by the survey. Right panel: Distribution of targets and associated tiling map for the tNCCSz survey. Color-coded by magnitude, the map highlights galaxies in a specific magnitude range for redshift analysis. The circular red outlines represent the 56 observational tiles, each corresponding to a unique pointing of the WIYN telescope to cover the declination range above $87^{\circ}$ and parts of the 86.6 to 87.0 range. The blued dashed circle represents the full width at half maximum of the primary beams of the TDA antennas at 1380 MHz ($z\sim 0.035$).
@@ II The Survey
To conduct an optical galaxy redshift survey, one must first identify the angular positions of the galaxies using a photometric survey and then perform spectroscopic observations. The photometry for our study was previously done in the North Celestial Cap Survey (NCCS Gorbikov and Brosch (2014)). The NCCS identifies approximately $4\times 10^{6}$ objects within $10^{\circ}$ of the NCP using V-R-I photometry with magnitude limits $m_{V}<20.3$, $m_{R}<21.0$, and $m_{I}<19.2$. Some of these objects are stars, and some galaxies, particularly at the magnitude limit, are likely beyond our primary interest in redshifts with $z<=0.07$. Thus, only a fraction of the NCCS survey is useful for our purposes. The catalog also includes a PESS (point-extended-source-separation) index indicating whether an object is point-like (PESS <=<= 1, most likely a star) or extended (PESS = 2 or 3, most likely a galaxy). Selecting spectroscopic targets involved cutting on declination and magnitude and excluding point-like objects.
@@ II.2 Observations
Between February 2020 and May 2023, we obtained 15 nights of gray and dark time on the WIYN 3.5 m telescope to use the Hydra multifiber positioner and bench spectrograph (Barden and Armandroff, 1995) to survey the NCC. We used the bench spectrograph with the blue-light optimized fibers and the 400@4.2 (WIYN Hydra and Bench spectrograph users manual - v5) grating. These choices resulted in wavelength coverage from below 3800 Å to above 7600 Å. However, our sensitivity below about 4600 Å was typically insufficient to identify the common absorption features in galaxy spectra, and the $\textrm{H}_{\alpha}$ emission line could be confused with prominent sky lines above 7200 Å.
@@ III.3.2 Semi-automatic Method
We chose to use two (for robustness) galaxy spectra templates with emission features and two with absorption features. Emission line galaxy spectra have been chosen from the SDSS DR5 spectral cross-correlation templates [^2] https://classic.sdss.org/dr5/algorithms/spectemplates/. Specifically, we used templates with IDs 24 and 25 in our processing. For absorption line galaxies, on the one hand, we used the NGC 4125 spectrum listed by Brown et al. (2014) in their representative subsample. We also used template no. 41 in the collection used in the AUTOZ code developed within the GAMA collaboration (Baldry et al., 2014)333code and data can be found in I. Baldry’s website: https://www.astro.ljmu.ac.uk/ ikb/research/. Finally, two redshift measurements have been derived using template ID 30 in the SDSS DR5 collection, corresponding to objects with broad emission lines (like QSOs or AGN). Prior to their use in redshift determination, all templates are pre-processed by fitting and subtracting a low order polynomial.
@@ IV.1.1 A posteriori check of target selection with Gaia
In order to estimate the fraction of proposed targets that might actually be stars, rather than galaxies, we matched the NCCS catalog with the Gaia DR3 catalog (Gaia Collaboration et al., 2023). As this catalog was released well after the beginning of our survey, this is only an a posteriori check. First, we looked at the distribution of the PESS flag for NCCS objects with $\delta\geq 86.6^{\circ}$ and $13<m_{V}<18.995$ - essentially the same as the criteria used for our target selection - shown on Figure 6. Within this sample, 1714 and 669 objects have PESS flag equal to 2 and 3, respectively. Of these, 1046 objects with PESS = 2, and 222 objects with PESS = 3, are closer than 2 arcsec from an object in Gaia DR3 (a “match” in the following). In contrast, more than 99.8% of NCCS objects with PESS $\leq$ 1 (point-like) have a match in Gaia DR3.
@@ IV.1.3 Comparison with the 2MASS Redshift Survey
The 2MASS redshift survey catalog (Huchra et al., 2012) contains a number of galaxies that overlap our survey area. As a means for testing the completeness of the tNCCSz, we searched the 2MASS catalog for galaxies with $\delta\geq 86.5^{\circ}$ and $m_{K}>13.5$. We found 26 such galaxies and
our tNCCSz survey obtained spectra for 14 of them. The breakdown of the reasons the remaining 12 were not in our sample is summarized in Table 2. The two targets that were “Not in the NCCS Catalog” were detected in V-band by NCCS but without corresponding measurements in R or I bands. These objects appear in the full NCCS catalog but not in the “cleaned” catalog that we used for target selection. Somewhat surprisingly, 10 galaxies were not included in our target list because the NCCS catalog PESS score indicated that they were more likely stars. This result appears to be inconsistent with our understanding that the NCCS catalog separation between stars and galaxies is $92\%$ accurate, based on a comparison of NCCS with SDSS (Gorbikov and Brosch, 2014).
@@ IV.2 Selection Function
Using the above efficiency functions, we are able to reproduce the magnitude distribution of our final list of galaxies with redshifts, the NCCS catalog itself, and the selected targets for this survey. To do so, we start from the galaxy absolute luminosity distribution function, and we take into account the absorption by the Galactic dust, with its variations over the $\sim 35\,\mathrm{deg}^{2}$ survey area (Figure 1, right panel). The galaxy luminosity distribution is modeled as a Schechter function (see Eq. 7) with the parameters given by Zucca et al. (2009). More precisely, we have used a characteristic magnitude $M^{*}=-20.5$, normalisation $\Phi^{*}=0.0063\mathrm{Mpc}^{-3}$ and a faint end slope $\alpha=-1.07$ as the luminosity function parameters. More details can be found in paragraph V.3.1.
@@ IV.3 Catalog Description
The tNCCSz redshift catalog contains a total of 23 columns and 1874 rows, with each row corresponding to an NCCS (Gorbikov and Brosch, 2014) source, with the associated redshift information extracted from the current survey’s observations. Eleven columns correspond to information in the original NCCS catalog, and an additional 12 columns contain redshift information derived from this survey. The full table in electronic form will be available from the the Centre de Données astronomiques de Strasbourg (CDS) [^5] CDS, Strasbourg astronomical Data Center: https://cds.unistra.fr.
@@ V Clustering analysis
One way to validate tNCCSz is to analyze the redshift space clustering of our galaxies and compare this with other surveys since the clustering statistics in different cosmological volumes are expected to be the same. Due to its small solid angle tNCCSz can add little new to our knowledge of galaxy clustering, but this is not why this survey was made. The comparison we make is with the Sloan Digital Sky Survey (SDSS). SDSS is chosen because of its similar depth, excellent control of systematics, and its very large area, which provides a large statistical sample. More specifically, we compare tNCCSz with mock tNCCSz catalogs derived from the SDSS spectroscopic catalog (Blanton et al., 2017). The procedure for generating these mocks is described in § V.1.